# FaceQSORT: a Multi-Face Tracking Method based on Biometric and Appearance Features

### v2

@@ 4.4.2 Selection of a dynamic $\lambda$ based on Face Quality
By setting the parameter $\lambda$ to a fixed value (e.g. $\lambda=0.1$), the weighting between biometric and optical features is constant for each image and each detected face, regardless of the biometric information present in the extracted face patch. For this reason, a more natural approach would be for $\lambda$ to reflect the available biometric information (i.e., $\lambda=1$ when full biometric information is available and $\lambda=0$ when no biometric information is available). The biometric information present in a face sample can be quantified by a face quality score [67, 68, 69]. For this reason, another option would be to set $\lambda$ dynamically according on a face quality score. This turns $\lambda$ into an $m\times n$ matrix with constant columns corresponding to the face quality scores of the $n$ detected faces. The multiplications $\lambda C_{bio}$ and $(1-\lambda)C_{app}$ are then element wise multiplications.
@@
Three different face quality models are exploited, i.e., CR-FIQA [67], MagFace [68] and FaceQgen [69]. In case of the CR-FIQA two different backbone model sizes are evaluated, i.e., CR-FIQA-S and CR-FIQA-L. To obtain a face quality score the detected face patch is fed into the respective model. The distribution of quality scores achieved for all detected faces (across all sequences) is illustrated in Figure 8. Quality values that are not restricted to $\mathbb{R}_{[0,1]}$ are normalized by dividing by the maximum observed value.
@@
Figure 9: Achieved AssA, IDF1 and IDSW with $\lambda$ dynamically set based on normalized (by the maximum value) face quality values compared to the best achieved AssA, IDF1 and IDSW based on a fixed $\lambda$ value. Recall: for AssA and IDF1, the higher the better, and for IDSW, the lower the better.
@@
In Figure 9, the results are compared between a fixed (best results table 2 and 3) and a dynamically set (based on face quality scores) $\lambda$. It can be seen that tracking performance is generally lower when $\lambda$ is dynamically set based on face quality scores. Only with CR-FIQA is the performance for some face descriptors as high as when $\lambda$ is set to a fixed value. In general, the performance based on CR-FIQA is consistently better than the other two face quality models evaluated. The better performance of CR-FIQA could be expected when looking at the distribution of face quality scores (Figure 8), as FaceQgen and MagFace only cover a limited range, while CR-FIQA covers almost the entire range from 0 to 1, which appears more natural. Thus, the reasonable approach of dynamically setting $\lambda$ based on available biometric information (face quality scores) is not working. A reason for this could be that the considered face quality scores do not reflect the actuall biometric information present.
4.4.3 Grid Search
To assess whether a higher AssA can be achieved with a dynamic $\lambda$ (which is different for each detected face) than with a fixed $\lambda$, a grid search is performed. In a grid search, all possible combinations of $\lambda$ are evaluated. For example, if 5 faces are detected per frame and 11 different $\lambda$ values are allowed (i.e. $\lambda\in[0.0,0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,1.0]$), this would result in $11^{5}$ different combinations for the frame under consideration. Since the association in the current frame depends on the past frames, the different combinations must be considered across all frames, i.e. $\prod_{i=1}^{N}11^{m^{i}}$ combinations, where $N$ is the number of frames and $m^{i}$ are the detected faces for the $i$-th frame. This is simply not computable. Thus, only a local grid search (between two consecutive frames) is performed. More precisely, the AssA is calculated after each $\lambda$ combination for the current frame. The combination that achieves the highest AssA is selected. If several combinations achieve the highest AssA, the combination that achieves the highest AssA first is selected.
(a) sequence ID 04
(b) sequence ID 03
@@ 4.4.3 Grid Search
Figure 10: Highest AssA achieved per frame in the grid search.
(a) sequence ID 04
(b) sequence ID 03
@@
Figure 11: Face patches that were assigned a $\lambda$ of 1.0 during the grid search.
@@
Since the evaluation of 11 different $\lambda$ values is still very time-consuming, a smaller number is considered. In addition, the grid search is only performed with the best face descriptor (in terms of AssPr), SFace, and only for the two sequences with the fewest frames (sequence ID 3 and 4). In Figure 10 (a), the highest AssA achieved for each frame of sequence 4 is illustrated. It can be seen that the AssA at frame 576 (complete sequence) is about 0.09 higher than the baseline (fixed $\lambda$) and exactly the same results are obtained regardless of whether $\lambda$ is selected from [1.0,0.9,0.7,0.5,0.3,0.1,0.0] or [1.0,0.1,0.0].
@@
Taking into account possible $\lambda$ values of [1.0,0.9,0.7,0.5,0.3,0.1,0.0], a $\lambda$ of 1.0 is selected for 2199/2239 detected faces in sequence 04. The $\lambda$ values for the remaining 40 detected faces are distributed as follows: 5, 22, 7, 4, 2 for the $\lambda$ values 0.0, 0.1, 0.3, 0.5 and 0.7, respectively. The vast majority of $\lambda$ values are set to 1.0 because the grid search conducted always selects the first $\lambda$ combination that achieves the highest AssA value, and the first combination tried is that all $\lambda$ values equal 1.0. In the version where $\lambda$ can be chosen from only three values (i.e. [0.1,0.0,1.0]), the first combination tried is that all $\lambda$ values are equal to 0.1, and the last combination is that all $\lambda$ values are equal to 1.0 (i.e. a $\lambda$ of 1.0 is only chosen if 0.1 and 0.0 do not achieve an equally high AssA). In this case, a $\lambda$ of 0.1 is selected for 2206/2239 detected faces. Compared to the baseline (i.e. fixed $\lambda$ of 0.1 for all recognized faces), a change of $\lambda$ for only 33 recognized faces yields the performance gain. In particular, for 24 detected faces, a $\lambda$ value of 1.0 is selected. A $\lambda$ value of 1.0 means that only face features are used. Therefore, it is expected that the corresponding face patches contain adequate biometric information. The 24 face patches where $\lambda$ is set to 1.0 are illustrated in Figure 11 (a). As can be seen, the biometric information is very limited.
The results for sequence 3 are similar. In Figure 10 (b), the highest AssA achieved for each frame is illustrated. If the possible $\lambda$ values are $[0.1,0.0,1.0,0.5]$, a value of 0.1 is selected for 2431/2466 detected faces. A $\lambda$ value of 0.0, 0.5 and 1.0 is selected for 6, 1 and 28 recognized faces, respectively. Again, it would be expected that with a selected $\lambda$ value of 1.0, there should be reasonable biometric information in the face patches. However, as can be seen in Figure 11 (b) the biometric information in all 28 face patches is limited.
@@
These observations indicate that linking $\lambda$ to the available biometric information does not lead to better performance.
@@ 4.4.4 Matching Score Distribution
Table 4: Presents the overlap (IoU) between the genuine and imposter scores as well as their mean ($\mu$) and standard deviation ($\sigma$). Feature models: (1) ArcFace, (2) DeepFace, (3) DeepID, (4) Dlib, (5) Facenet, (6) Facenet512, (7) OpenFace, (8) SFace, (9) SeNet50, (10) VGG-Face, (11) ArcFaceII and (12) Appearance Features.
@@
A multi-face tracker can be compared with a biometric system (e.g., a set of query images (detected faces) are matched against a database of enrolled identities (active tracks)). Matching is based on a similarity score, and in a perfect system, the genuine scores do not overlap with the impostor scores. In the context of FaceQSORT, the genuine scores represent the cosine similarity that result when a detected face is compared with the corresponding stored faces (active tracks). The imposter scores, on the other hand, represent the cosine similarities that are achieved when a detected face is compared with all other stored faces (tracks). Table 4 shows the mean and the standard deviation of the genuine and the imposter score distribution as well as their overlap (IoU). Considering the IoU, the most distinct face descriptor (lowest IoU) is SFace. Other face descriptors with a low IoU are Facenet512, VGG-Face, ArcFaceII and Dlib. These face descriptors (except Dlib) are also the face descriptors achieving the highest AssPr (see Figure 7). However, based on the IoU results, it is surprising that such low AssPr values are obtained with the Dlib face descriptor. A reason for the low AssPr values could be that, with a mean value of 0.07, the imposter scores are well below the threshold $\theta=0.2$ (applied to reject uncertain (imposter) matches). Thus, a closer look is taken at the selection of the threshold value $\theta$.
@@ 4.4.5 Selection of the Threshold $\theta$
Table 5: Best achieved AssA, IDF1 and IDSW for different selections of $\theta$ and $\lambda$ combinations. Face descriptors: (1) ArcFace, (2) DeepFace, (3) DeepID, (4) Dlib, (5) Facenet, (6) Facenet512, (7) OpenFace, (8) SFace, (9) SeNet50, (10) VGG-Face and (11) ArcFaceII.
@@
Figure 12: AssRe over AssPr for different $\theta$ and $\lambda$ combinations and SFace features. The $\theta$ values increase from left to right and the contour lines represent the AssA, where the area shaded in red indicates a lower AssA than for the best achieved AssA when only appearance features are used.
@@
For both datasets, the genuine and impostor scores obtained on the basis of the Dlib features are close to zero (see Table 4). Thus, with a threshold $\theta=0.2$ possible imposter matches (inter-person identity switches) are not rejected. This explains the low AssPr value when only Dlib features are used for the association ($\lambda=1.0$), although the genuine and imposter distributions are relatively distinct (an IoU value of 0.07 and 0.08). If $\theta$ is set to 0.025, an AssPr of 0.9831 and 0.9938 is achieved for the PLUSFiaQ and ChokePoint datasets respectively. This shows that, in addition to $\lambda$, $\theta$ is an important parameter as well. In Table 5, the best achieved AssA, IDF1 and IDSW values for different $\theta$ and $\lambda$ combinations are shown. It can be seen that the initial choice of $\theta=0.2$ was already pretty good for several face descriptors. Overall, the best results are achieved with SFace, which also provides the most distinctive, genuine and imposter score distributions (IoU of 0.04). Again, the $\lambda$ values with which the best results are achieved are significantly higher for the ChokePoint dataset (probably less occlusions and lateral faces benefit biometric features). Although very low $\theta$ values (e.g. 0.0125, 0.025, 0.05, etc.) were evaluated, no combination of Dlib and appearance features achieved higher tracking performance than with appearance feature only (i.e. $\lambda=0.0$).
@@
When considering the threshold $\theta$ the expectation is that the AssPr is very high for low $\theta$ values (reject most imposter matches) and it decreases when $\theta$ increases. However, a low value of $\theta$ can lead to a lower AssRe (more intra-person trID switches) since also genuine matches are rejected. In Figure 12 the trade-off between AssRe and AssPr compared to AssA (contour lines) for the best face descriptor SFace is shown for different $\lambda$ and $\theta$ combinations. The $\theta$ values decrease from left to right (i.e., 0.9, 0.8, 0.7, 0.6, 0.5, 0.4, 0.3, 0.2, 0.1, 0.05 and 0.025) and the red lines show the AssRe and AssPr for the best AssA achieved with appearance features only (baseline). The area shaded in red represents the area where the AssA is lower than the baseline. A good trade-off (relatively constant high AssPr while reducing intra-person IDSWs (AssRe)) when increasing the threshold $\theta$ can be achieved for higher $\lambda$ values (higher contribution of biometric features).
4.4.6 Selection of the Similarity Metric
Table 6: L2: Best achieved AssA, IDF1 and IDSW for different selections of $\theta$ and $\lambda$ combinations. Face descriptors: (1) ArcFace, (2) DeepFace, (3) DeepID, (4) Dlib, (5) Facenet, (6) Facenet512, (7) OpenFace, (8) SFace, (9) SeNet50, (10) VGG-Face and (11) ArcFaceII.
PLUSFiaQ ChokePoint
[…]
(1) 0.55/0.1 0.6287 0.55/0.1 0.7325 0.85/0.1 0.0080 0.55/0.3 0.8365 0.55/0.3 0.8801 1.15/0.3 0.0051
(2) 0.6/0.4 0.6398 0.6/0.4 0.7375 0.8/0.1 0.0080 0.7/0.6 0.8410 0.7/0.5 0.8839 0.9/0.9 0.0054
(3) 0.49/0.1 0.6215 0.49/0.1 0.7264 0.56/0.1 0.0086 0.6/0.4 0.8371 0.6/0.4 0.8803 0.8/0.8 0.0057
(4) 0.4/0.6 0.6444 0.4/0.6 0.7338 0.56/0.0 0.0091 0.4/0.6 0.8399 0.4/0.6 0.8793 0.4/0.8 0.0078
(5) 0.55/0.1 0.6395 0.55/0.1 0.7389 0.85/0.2 0.0080 0.55/0.3 0.8310 0.55/0.3 0.8723 1.15/0.4 0.0051
(6) 0.7/0.4 0.6762 0.7/0.4 0.7560 0.85/0.3 0.0079 0.85/0.5 0.8539 0.85/0.5 0.8882 1.15/0.4 0.0051
(7) 0.6/0.4 0.6398 0.7/0.6 0.7382 0.8/0.2 0.0080 0.6/0.3 0.8272 0.6/0.3 0.8731 0.9/0.6 0.0055
(8) 0.85/0.4 0.6672 0.85/0.4 0.7636 1.15/0.7 0.0074 1.15/0.9 0.8800 1.0/0.6 0.9010 1.3/0.3 0.0047
(9) 0.49/0.3 0.6376 0.49/0.3 0.7353 0.56/0.4 0.0081 0.6/0.4 0.8409 0.6/0.4 0.8781 0.8/0.6 0.0050
(10) 0.7/0.4 0.6565 0.7/0.4 0.7469 0.85/0.1 0.0079 0.7/0.6 0.8562 0.85/0.6 0.8872 1.15/0.9 0.0051
(11) 0.7/0.3 0.6613 0.7/0.3 0.7520 1.0/0.6 0.0077 1.0/0.8 0.8574 1.0/0.8 0.8925 1.0/0.7 0.0051
$\varnothing$ 0.60/0.32 0.6466 0.61/0.34 0.7419 0.80/0.25 0.0081 0.70/0.52 0.8455 0.70/0.48 0.8833 0.97/0.61 0.0054
Table 7: L2: Presents the overlap (IoU) between the genuine and imposter scores as well as their mean ($\mu$) and standard deviation ($\sigma$). Feature models: (1) ArcFace, (2) DeepFace, (3) DeepID, (4) Dlib, (5) Facenet, (6) Facenet512, (7) OpenFace, (8) SFace, (9) SeNet50, (10) VGG-Face, (11) ArcFaceII and (12) Appearance Features.
PLUSFiaQ ChokePoint
genuine imposter genuine imposter
IoU $\mu/\sigma$ $\mu/\sigma$ IoU $\mu/\sigma$ $\mu/\sigma$
(1) 0.26 0.45/0.26 0.95/0.37 0.20 0.48/0.27 1.01/0.35
(2) 0.14 0.41/0.15 0.70/0.11 0.12 0.43/0.15 0.72/0.11
(3) 0.20 0.18/0.12 0.41/0.14 0.17 0.20/0.13 0.43/0.13
(4) 0.07 0.19/0.06 0.37/0.06 0.08 0.21/0.06 0.38/0.07
(5) 0.18 0.43/0.22 0.97/0.28 0.26 0.44/0.25 0.90/0.32
(6) 0.07 0.52/0.23 1.22/0.17 0.09 0.57/0.23 1.19/0.20
(7) 0.18 0.43/0.16 0.77/0.17 0.18 0.45/0.16 0.81/0.20
(8) 0.04 0.71/0.22 1.31/0.10 0.04 0.74/0.20 1.34/0.09
(9) 0.18 0.19/0.10 0.41/0.13 0.19 0.28/0.15 0.67/0.23
(10) 0.07 0.52/0.19 1.09/0.14 0.08 0.60/0.18 1.14/0.17
(11) 0.08 0.61/0.22 1.23/0.15 0.05 0.65/0.19 1.27/0.13
(12) 0.16 0.20/0.11 0.70/0.29 0.18 0.23/0.13 0.67/0.28
Another factor that can influence the tracking performance is the selection of the similarity metric (used to compue $C_{i,j}^{app/bio}$). The best results achieved when using Euclidean distance instead of cosine similarity are shown in Table 6. Again, the overall best results are achieved with SFace features, which are partly better than when using cosine similarity. However, the average performance across all face descriptors assessed is slightly lower. Another noticeable difference is that the $\theta$ and $\lambda$ values with which the best results are achieved are significantly higher. The mean and standard deviation of the genuine and imposter score distributions in terms of the Euclidean distance is shown in Table 7. It can be seen that they are clearly higher than in Table 4 (cosine similarity). The IoUs, however, are identical. The identical IoUs could be due to the fact that the features are normalized before calculating the Euclidean distance. Thus, when using the Euclidean distance, the scores are more or less scaled differently. For the differently scaled scores, better performance can also be achieved with Dlib features using a combination of biometric and appearance features (i.e. $\lambda=0.6$).
4.4.7 Compare Tracker

### v3

@@ 6.1.2 Selection of a dynamic $\lambda$ based on Face Quality
By setting the parameter $\lambda$ to a fixed value (e.g., $\lambda=0.1$), the weighting between biometric and appearance features is constant for each image and each detected face, regardless of the biometric information present in the extracted face patch. For this reason, a more natural approach would be for $\lambda$ to reflect the available biometric information (i.e., $\lambda=1$ when full biometric information is available and $\lambda=0$ when no biometric information is available). The biometric information present in a face sample can be quantified by a face quality score [61, 62, 63]. For this reason, another option would be to set $\lambda$ dynamically according to a face quality score. This turns $\lambda$ into an $|\Phi|\times|\Psi|$ matrix with constant columns corresponding to the face quality scores of the detected faces $\Psi$. The multiplications $\lambda C_{bio}$ and $(1-\lambda)C_{app}$ are then element wise multiplications.
@@
Three different face quality models are exploited, i.e. CR-FIQA [61], MagFace [62] and FaceQgen [63]. In case of the CR-FIQA two different backbone model sizes are evaluated, i.e. CR-FIQA-S and CR-FIQA-L. To obtain a face quality score, the detected face patch is fed into the respective model. The distribution of quality scores achieved for all detected faces (across all sequences) is illustrated in Figure 8. Quality values that are not restricted to $\mathbb{R}_{[0,1]}$ are normalized by dividing by the maximum observed value.
@@
Figure 9: Achieved AssA, IDF1 and IDSW with $\lambda$ dynamically set based on normalized (by the maximum value) face quality scores compared to the best achieved AssA, IDF1 and IDSW based on a fixed $\lambda$ value. Recall: for AssA and IDF1, the higher the better, and for IDSW, the lower the better.
@@
In Figure 9, the results are compared between a fixed (best results table 2 and 3) and a dynamically set (based on face quality scores) $\lambda$. It can be seen that tracking performance is generally lower when $\lambda$ is dynamically set based on face quality scores. Only with CR-FIQA the performance for some face recognition models is as high as when $\lambda$ is set to a fixed value. In general, the performance based on CR-FIQA is consistently better than the other two face quality models evaluated. However, the reasonable approach of dynamically setting $\lambda$ based on available biometric information (face quality scores) is not working. One reason for this could be that face quality metrics are designed to predict face recognition performance, which does not directly translate to multi-face tracking.
6.1.3 Grid search dynamic $\lambda$
@@
Figure 10: Highest AssA achieved per frame in the grid search (sequence 4).
To assess whether a higher AssA can be achieved with a dynamic $\lambda$ (which is different for each detected face) than with a fixed $\lambda$, a grid search is performed. In a grid search, all possible combinations of $\lambda$ are evaluated. For example, if 5 faces are detected per frame and 11 different $\lambda$ values are allowed (i.e., $\lambda\in[0.0,0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,1.0]$), this would result in $11^{5}$ different combinations for the frame under consideration. Since the association in the current frame depends on the past frames, the different combinations must be considered across all frames, i.e. $\prod_{i=1}^{N}11^{m^{i}}$ combinations, where $N$ is the number of frames and $m^{i}$ are the detected faces for the $i$-th frame. This is simply not computable. Thus, only a local grid search (between two consecutive frames) is performed. More precisely, the AssA is calculated after each $\lambda$ combination for the current frame. The combination that achieves the highest AssA is selected. If several combinations achieve the highest AssA, the combination that achieves the highest AssA first is selected.
@@
Figure 11: Face patches that were assigned a $\lambda$ of 1.0 during the grid search with sequence 4. Since only biometric features are used for $\lambda=1.0$, it is expected that the corresponding face patches contain rich biometric information, which is not the case.
@@
Since the evaluation of 11 different $\lambda$ values is still very time-consuming, a smaller number is considered. In addition, the grid search is only performed with the best face recognition model (in terms of AssPr), BIO-SF01, and only for the two sequences with the fewest frames (sequence ID 3 and 4). In Figure 10, the highest AssA achieved for each frame of sequence 4 is illustrated. It can be seen that the AssA at frame 576 (complete sequence) is about 0.09 higher than the baseline (fixed $\lambda$) and exactly the same results are obtained regardless of whether $\lambda$ is selected from {1.0,0.9,0.7,0.5,0.3,0.1,0.0} or from {0.1,0.0,1.0}.
@@
Taking into account possible $\lambda$ values of {1.0,0.9,0.7,0.5,0.3,0.1,0.0}, a $\lambda$ of 1.0 is selected for 2199/2239 detected faces in sequence 4. The $\lambda$ values for the remaining 40 detected faces are distributed as follows: 5, 22, 7, 4, 2 for the $\lambda$ values 0.0, 0.1, 0.3, 0.5 and 0.7, respectively. The vast majority of $\lambda$ values are set to 1.0 because the grid search conducted always selects the first $\lambda$ combination that achieves the highest AssA value, and the first combination tested is that all $\lambda$ values equal 1.0. In the version where $\lambda$ can be chosen from only three values (i.e., {0.1,0.0,1.0}), the first combination tested is that all $\lambda$ values are equal to 0.1, and the last combination is that all $\lambda$ values are equal to 1.0 (i.e., a $\lambda$ of 1.0 is only chosen if 0.1 and 0.0 do not achieve an equally high AssA). In this case, a $\lambda$ of 0.1 is selected for 2206/2239 detected faces. Compared to the baseline (i.e., fixed $\lambda$ of 0.1 for all detected faces), a change of $\lambda$ for only 33 detected faces yields the performance gain. In particular, for 24 detected faces, a $\lambda$ value of 1.0 is selected. A $\lambda$ value of 1.0 means that only biometric features are used. Therefore, it is expected that the corresponding face patches contain adequate biometric information. The 24 face patches where $\lambda$ is set to 1.0 are illustrated in Figure 11. As can be seen, the biometric information is very limited.
@@
The results for sequence 3 are similar. These observations indicate that linking $\lambda$ to the available biometric information does not lead to better performance.
@@ 6.1.4 Matching score distribution
Table 4: Presents the overlap (IoU) between the genuine and imposter scores as well as their mean ($\mu$) and standard deviation ($\sigma$). Feature models: (1) BIO-AF01, (2) BIO-AF02, (3) BIO-DF01, (4) BIO-DI01, (5) BIO-DL01, (6) BIO-FN01, (7) BIO-FN02, (8) BIO-OF01, (9) BIO-SF01, (10) BIO-SN01, (11) BIO-VG01 and (12) Appearance Features.
@@
A multi-face tracker can be compared with a biometric system (e.g., a set of query images (detected faces) are matched against a database of enrolled identities (active tracks)). Matching is based on a similarity score, and in a perfect system, the genuine scores do not overlap with the impostor scores. In the context of FaceQSORT, the genuine scores represent the cosine similarity that result when a detected face is compared with the corresponding stored faces (active tracks). The imposter scores, on the other hand, represent the cosine similarities that are achieved when a detected face is compared with all other stored faces (tracks). Table 4 shows the mean and the standard deviation of the genuine and the imposter score distribution, as well as their overlap (IoU). Considering the IoU, the most distinct face recognition model (lowest IoU) is BIO-SF01. Other face recognition models with a low IoU are BIO-FN02, BIO-VG01, BIO-AF02 and BIO-DL01. These face recognition models (except BIO-DL01) are also the models achieving the highest AssPr (see Figure 7). However, based on the IoU results, it is surprising that such low AssPr values are obtained with the Dlib face recognition model. A reason for the low AssPr values could be that, with a mean value of 0.07, the imposter scores are well below the threshold $\theta=0.2$ (applied to reject uncertain (imposter) matches). Thus, a closer look is taken at the selection of the threshold value $\theta$.
@@ 6.1.5 Selection of the threshold $\theta$
Table 5: Best achieved AssA, IDF1 and IDSW for different selections of $\theta$ and $\lambda$ combinations. Face recognition models: (1) BIO-AF01, (2) BIO-AF02, (3) BIO-DF01, (4) BIO-DI01, (5) BIO-DL01, (6) BIO-FN01, (7) BIO-FN02, (8) BIO-OF01, (9) BIO-SF01, (10) BIO-SF01 and (11) BIO-VG01.
@@
Figure 12: AssRe over AssPr for different $\theta$ and $\lambda$ combinations and BIO-SF01 features. The $\theta$ values increase from left to right and the red lines represent the baseline (AssRe and AssPr of the best achieved AssA, if only appearance features are used).
@@
For both datasets, the genuine and impostor scores obtained on the basis of BIO-DL01 features are close to zero (see Table 4). Thus, with a threshold $\theta=0.2$ possible imposter matches (inter-person trID switches) are not rejected. This explains the low AssPr value when only BIO-DL01 features are used for the association (see ‘1) Selection of a fixed $\lambda$’ of subsection 6.1), although the genuine and imposter distributions are relatively distinct (an IoU value of 0.07 and 0.08). If $\theta$ is set to 0.025, an AssPr of 0.9831 and 0.9938 is achieved for the PLUSFiaQ and ChokePoint datasets, respectively. This shows that, in addition to $\lambda$, $\theta$ is an important parameter as well. In Table 5, the best achieved AssA, IDF1 and IDSW values for different $\theta$ and $\lambda$ combinations are shown. It can be seen that the initial choice of $\theta=0.2$ was already pretty good for several face recognition models. Overall, the best results are achieved with BIO-SF01 features, which also provides the most distinctive genuine and imposter score distributions (IoU of 0.04). Again, the $\lambda$ values that yield the best results are significantly higher for the ChokePoint dataset (probably fewer occlusions and lateral faces benefit the usage of biometric features). Although very low $\theta$ values (e.g. 0.025, 0.05, etc.) are tested, a combination of BIO-DL01 and appearance features performs worst overall.
@@
When considering the threshold $\theta$ the expectation is that the AssPr is very high for low $\theta$ values (reject most imposter matches) and it decreases when $\theta$ increases. However, a low value of $\theta$ can lead to a lower AssRe (more intra-person trID switches) since also genuine matches are rejected. In Figure 12 the trade-off between AssRe and AssPr for the best face recognition model BIO-SF01 is shown for different $\lambda$ and $\theta$ combinations. The $\theta$ values decrease from left to right (i.e., 0.9, 0.8, 0.7, 0.6, 0.5, 0.4, 0.3, 0.2, 0.1, 0.05 and 0.025) and the red lines show the AssRe and AssPr for the best AssA achieved with appearance features only (baseline). A good trade-off (relatively constant high AssPr while reducing intra-person trID switches (AssRe)) when increasing the threshold $\theta$ can be achieved for higher $\lambda$ values (higher contribution of biometric features).
6.2 Tracker Comparison